OpenStudy (anonymous):

I got a total of 777 people.... 47% of them want their windows was he by hiring someone to do it i got 365 people as an answer... but i have a plus or minus error of 3.52% can you help me get my answer? is it 3.52 of 777?

OpenStudy (anonymous):

i get 27 as a whole number since were talking about people.

OpenStudy (anonymous):

3.52% =\dfrac {3.52}{100}= 0.0352 So that 3.52% of 777 means 0.0352*777

OpenStudy (anonymous):

yeah so is my answer 365 +27.... and or 365-27??

OpenStudy (anonymous):

\(\pm\) not just -

OpenStudy (anonymous):

so leave it as 365 +-27

OpenStudy (anonymous):

If you have to round, you must let the whole number (including decimal) to calculate, then round. For example : 365.19 - 27.35=367.84 then round to 368 365.19+ 27.35= 392.54, then round to 393 If you don't have to, you let the whole thing as it is \(365.19\pm 27.35\)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ok